Fogerty Company makes two products—titanium Hubs and Sprockets. Data regarding the two products follow:

Direct
Labor-Hours per Unit
Annual
Production
Hubs 0.70 13,000 units
Sprockets 0.30 54,000 units

Additional information about the company follows:

  1. Hubs require $26 in direct materials per unit, and Sprockets require $11.

  2. The direct labor wage rate is $17 per hour.

  3. Hubs require special equipment and are more complex to manufacture than Sprockets.

  4. The ABC system has the following activity cost pools:

Estimated Activity
Activity Cost Pool (Activity Measure) Overhead Cost Hubs Sprockets Total
Machine setups (number of setups) $ 18,900 105 84 189
Special processing (machine-hours) $ 253,000 4,600 0 4,600
General factory (organization-sustaining) $ 106,300 NA NA NA

Required:

1. Compute the activity rate for each activity cost pool.

2. Determine the unit product cost of each product according to the ABC system.

Determine the unit product cost of each product according to the ABC system. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Hubs Sprockets
Direct materials $26.00 $11.00
Direct labor 11.90 5.10
Overhead
Unit cost $37.90 $16.10

I'm trying to find the overhead cost. Needs some help
Answer #1

1) Overhead rate :

Machine setup = 18900/189 = 100 per setup

Special processing = 253000/4600 = 55 per machine hour

Determine the unit product cost of each product according to the ABC system. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Hubs Sprockets
Direct materials $26.00 $11.00
Direct labor 11.90 5.10
Overhead 263500/13000 = 20.27 8400/54000 = 0.16
Unit cost $58.17 $16.26